HETAV PATEL

Grade 11,

In acidic medium, equivalent weight of K 2 Cr 2 O 7 (Molar weight= M) is M/3 M/4 M/6 M/2 Do explain your answer.

In acidic medium, equivalent weight of K2Cr2O7 (Molar weight= M) is
  1. M/3
  2. M/4
  3. M/6
  4. M/2
Do explain your answer.

Grade:12th pass

3 Answers

Rachna Verma
4 years ago
The correct answer is M/3 because the oxidation state of Cr changes from 6 to 3 in presence of acidic medium
Vikas TU
4 years ago
K2Cr2O7 in solution gives following ions -
K2Cr2O7 + 14H+ + 6e- → 2K+ + 2Cr3+ + 7H2O
 
Here, 6 e- participate in the reaction, hence chemical equivalent is 6.
 
Equivalent weight is calculated by -
Z = W / E
Z = 294 / 6
=> M/6 
 
Arun
4 years ago
 

molecular weight of pottasium dichromate = M , X factor in acidic medium  = 6

equivalent weight = molecular weight / X factor 

                         = M /6 (option C)